OpenStudy (anonymous):

@ganeshie8 plz help.

OpenStudy (anonymous):

will you help me @ganeshie8

OpenStudy (anonymous):

if your not able to ill ask someone else?

OpenStudy (anonymous):

you got a question dear?

OpenStudy (anonymous):

@satellite73 yes could you help plz?

OpenStudy (anonymous):

maybe hope it is not as epic as the last one took like two hours maybe

OpenStudy (anonymous):

2. Holly is studying the growth of a particular species of plant. She writes the following equation to show the height of the plant f(n), in cm, after n days:\[f(n)=10(1.02)^{n}\]

OpenStudy (anonymous):

part 1 Starting from 0 days, plug x-values into our function to find out how many days it takes for the plant to grow to 11.04 cm. Write a compound inequality to show the reasonable domain of the plant growth:

OpenStudy (anonymous):

part 2 What is the y-intercept of our function? What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (anonymous):

there is a part 3 but i dont want to bother you much.

OpenStudy (anonymous):

So lets start @satellite73

OpenStudy (anonymous):

the faster the better!

OpenStudy (anonymous):

\[f(n)=10(1.02)^{n}\]

OpenStudy (anonymous):

try \[10(1.02)^2\] and \[10(1.02)^3\] etc until you get \(11.04\)

OpenStudy (anonymous):

of course you wont give me the answer, right, you know how I work.

OpenStudy (anonymous):

use a calculator should take 30 seconds at most

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got it on the second try

OpenStudy (anonymous):

giving u the answer would not make anyone learn

OpenStudy (anonymous):

i don't even know it xD

OpenStudy (anonymous):

i can't, i just get close to it in the calculator, 11.26.. that stuff

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=10%281.02%29^5

OpenStudy (anonymous):

looks like 5 does it

OpenStudy (anonymous):

@Survivor people do that, so thats why i tell them not to.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

yeah, i did, to the 6,7,8 powers, XD

OpenStudy (anonymous):

i forgot 5! So since we know what would get 11.04, what am i supposed to do afterwards?

OpenStudy (anonymous):

have a beer?

OpenStudy (anonymous):

go on to part b?

OpenStudy (anonymous):

Write a compound inequality to show the reasonable domain of the plant growth, this part?

OpenStudy (anonymous):

oh this again i guess you can't go back in time so i would say \(x\geq 0\)

OpenStudy (anonymous):

have a beer?! what did you try to say?

OpenStudy (anonymous):

try to say that we were done with that part, but i guess i didn't see the second part

OpenStudy (anonymous):

how long do you want your plant to live? some plants live a long time

OpenStudy (anonymous):

i get to put how long?

OpenStudy (anonymous):

it asks for the domain right? then number of days

OpenStudy (anonymous):

if so for 6 years

OpenStudy (anonymous):

so we can just say \(n\geq 0\) and let the plant live forever

OpenStudy (anonymous):

or we can say \[0\leq n \leq 2160\] and let the plant live for 6 years (which has 2160 days)

OpenStudy (anonymous):

whoo! a little lower then, 3 years

OpenStudy (anonymous):

what difference does it make? let the poor plant live now what is part b i forget

OpenStudy (anonymous):

Write a compound inequality to show the reasonable domain of the plant growth

OpenStudy (anonymous):

ok lets let it live 6 years and compute \[10(1.02)^{2160}\]

OpenStudy (anonymous):

so put that in the calculator?

OpenStudy (anonymous):

yeah that is too big lets let the sucker live for only one year

OpenStudy (anonymous):

yeah,pllleaase!

OpenStudy (anonymous):

\[10(1.02)^{365}\]

OpenStudy (anonymous):

so it would equal 13,774.08291966?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=10%281.02%29^%28365%29 pretty tall plant

OpenStudy (anonymous):

round

OpenStudy (anonymous):

to say \(13774\)

OpenStudy (anonymous):

so what would be our 'compound inequality'

OpenStudy (anonymous):

oh ok, round

OpenStudy (anonymous):

\[10\leq y \leq 13774\]

OpenStudy (anonymous):

what do we do with that number.

OpenStudy (anonymous):

nothing move on to part b already

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

what is b?

OpenStudy (anonymous):

just wait

OpenStudy (anonymous):

lol i am waiting....

OpenStudy (anonymous):

XD ok, so um... where in the world am I supposed to know that o_O

OpenStudy (anonymous):

i guess i have to scroll up and find d

OpenStudy (anonymous):

What is the y-intercept of our function? What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (anonymous):

part 2 What is the y-intercept of our function? What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (anonymous):

the y intercept is what you get when \(n=0\) in other words \(10\)

OpenStudy (anonymous):

it represents how tall the plant was when you started counting

OpenStudy (anonymous):

so it would be 1.02 or 10.2??

OpenStudy (anonymous):

it would be neither \[f(0)=10(1.02)^0=10\times 1=10\]

OpenStudy (anonymous):

oh ok, 10, gotcha!

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

What is the y-intercept of our function? What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (anonymous):

the y intercept is 10

OpenStudy (anonymous):

it represents how tall the plant was when you started counting

OpenStudy (anonymous):

now we need to know what the graph of the function represents.

OpenStudy (anonymous):

apparently it represents the growth of the plant

OpenStudy (anonymous):

that's what the graph of the function represents then?

OpenStudy (anonymous):

this is what it looks like for the first 20 days http://www.wolframalpha.com/input/?i=10%281.02%29^x+domain+0..20

OpenStudy (anonymous):

i was about to report abuse on you by accident. =3

OpenStudy (anonymous):

what would it be then,this, how tall the plant was when you started counting it represents the growth of the plant o_O

OpenStudy (anonymous):

yikes the y intercept is how tall when you started counting the graph represents the growth of the plant plotted vs time

OpenStudy (anonymous):

i should say the height of the plant vs time

OpenStudy (anonymous):

oh ok, understandable

OpenStudy (anonymous):

What is the average rate of change formula for exponential growth functions? last promise!!

OpenStudy (anonymous):

average rate of change over what time period?

OpenStudy (anonymous):

what do you mean over what time period?

OpenStudy (anonymous):

it asks for the average rate of change the rate of change is it increases by 2% per day, but the average rate depends on what time interval you are talking about what exactly does the question ask?

OpenStudy (anonymous):

What is the average rate of change formula for exponential growth functions? <-THIS EXACTLY

OpenStudy (anonymous):

2% per day

OpenStudy (anonymous):

So that would be it, 2% per day?

OpenStudy (anonymous):

yes because you are multiplying each day by \(1.02\)

OpenStudy (anonymous):

so the answer to, What is the average rate of change formula for exponential growth functions? would be 2% per day, just want to make sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

but sorry i forgot there was a part connected to this,Plug and chug n= 1 to n = 5 into your formula above.

OpenStudy (anonymous):

yw now you owe me another picture

OpenStudy (anonymous):

oh, let me see, i will look what i have

OpenStudy (anonymous):

ok you can do that use a calculator put in \[10\times(1.02)^1\\ 10\times (1.02)^2\\ 10\times (1.02)^3\\ 10\times (1.02)^4\\ 10\times (1.02)^5\]